OpenStudy (anonymous):

Calculate derivative of \[g(x)=cos(x)exp(x^{2})\]

geerky42 (geerky42):

Apply product rule. You take derivative of \(\cos x\), leave \(\exp(x^2)\) untouched, then you add it by derivative of \(\exp(x^2)\) and leave \(\cos x\) untouched. While taking derivative of \(\exp(x^2)\), apply chain rule there.

OpenStudy (anonymous):

chain rule was \[f(g(x)) \rightarrow f^{'}(g(x)).g^{'}(x)\] right ?

OpenStudy (turingtest):

yes

geerky42 (geerky42):

So you figured it out?

OpenStudy (anonymous):

Is second step \[f^{'}(co(x)exp(x^{2}).(-sin(x).exp(2x))\] ?

geerky42 (geerky42):

No, it's \(-\sin(x)\cdot\exp(x^2) + \cos(x)\cdot2x\exp(x^2)\)

OpenStudy (anonymous):

ok then i need to use this rule right ? \[fg=fg^{'}+f^{'}g\]

geerky42 (geerky42):

Derivative of \(\cos (x) \) is \(-\sin(x)\), and derivative of \(\exp(x^2)\) is \(2x\exp(x^2)\) So \(f'(x)\) would be \((\cos(x))'\cdot\exp(x^2)+\cos(x)\cdot(\exp(x^2))'\\~\\=−\sin(x)⋅\exp(x^2)+\cos(x)⋅2x\exp(x^2)\)

geerky42 (geerky42):

Yes, that rule.

geerky42 (geerky42):

Refresh page if you see question marks, lol

OpenStudy (anonymous):

ok i got it but only my problem is how to figure out when we which rule apply

OpenStudy (turingtest):

basically, when taking the derivative of the product of two functions, use the product rule\[D_x[f(x)g(x)]=f'(x)g(x)+f(x)g'(x)\]when taking the derivative of a composite function use the chain rule\[D_x(f[g(x)])=f'[g(x)]g'(x)\]

OpenStudy (anonymous):

ok TuringTest now its better to understand thank you and geerky both

OpenStudy (anonymous):

Write the binary number \[11,101\overline{110}\] as decimal fraction and explain it with help of geometric series.

OpenStudy (turingtest):

please post each question in a separate thread